OpenStudy (anonymous):

I need help understanding how to do a problem using linear approximations or differentials.

OpenStudy (anonymous):

post a problem

OpenStudy (anonymous):

\[\sqrt[3]{1001}\]

OpenStudy (anonymous):

y= x^1/3 y(1000)=10 \[y(x_1+\Delta x)=y' \Delta x + y(x)\] \[y'=\frac{1}{3}x^{-2/3}\] \[y(1001)=\frac{1}{3}1000^{-2/3}*(-1)+10\]

OpenStudy (anonymous):

Please disregard the last line

OpenStudy (anonymous):

\[y(1001)=\frac{1}{3}*1000^{-2/3}*(1)+10\]

OpenStudy (anonymous):

y(1001)=10.003

OpenStudy (anonymous):

any question?

OpenStudy (anonymous):

im writing it out step by step trying to understand it if i get stuck ill let you know. thanks!

OpenStudy (anonymous):

ok. Steps to these problem 1) find number close by that returns easy answers in our case it was 1000 since 1000^1/3 was just 10 plus it is close 1001 2) Write your general function ; y=x^1/3 in our case and differentiate 3) write this \[y(x_1+\Delta x)=y' \Delta x + y(x)\] \[\Delta x=+1 \] since 1001 is 1 bigger than 1000

OpenStudy (anonymous):

ok got it, is it the same as using y= f(a) + f'(a)(x-a)

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

oh okay could u explain using those terms please?

OpenStudy (anonymous):

remember y=mx+b ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this is actually the same m= f'(a) x=(x-a) b= f(a)

OpenStudy (anonymous):

ohh wow i see

OpenStudy (anonymous):

y= f(a) + f'(a)(x-a) a=1000 f(a)=10 \[f'(x)=\frac{1}{3}x^{-2/3}\] plug in 1000 f'(a)=\[\frac{1}{3}1000^{-2/3}\] x=1001 a=1000 (x-a)=1

OpenStudy (anonymous):

so its y=10 + 1/3 1000^-2/3(1)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay sec im gonna solve.

OpenStudy (anonymous):

okay i got 10.0033333333 which is wat u got thanks!

OpenStudy (anonymous):

Calc I?

OpenStudy (anonymous):

umm i think so.